OpenStudy (anonymous):

M(-2,-3) is the midpoint of RS. If S has coordinates (3,6) find the coordinates of R A) (-7,15) B) -7,-12) C) (8,-12) D) (8,15)

OpenStudy (anonymous):

( Sx + Rx ) / 2 = Mx (Sy+Ry)/2 = My Mx = -2 Sx + Rx = -4 Sx = 3 3 +Rx = -4 Rx = -7 My = -3 Sy+Ry = -6 Sy = 6 6 + Ry = -6 Ry = -12 R(-7,-12)